OpenStudy (emsxx3_):

need help quickly will fan and medal a circle has its center (-7' 5) and a radius of 3 units what is the equation of the circle

OpenStudy (emsxx3_):

@Jadeishere

OpenStudy (emsxx3_):

@agent0smith

OpenStudy (jadeishere):

idk sorry...

OpenStudy (emsxx3_):

please help quick @agent0smith

OpenStudy (agent0smith):

http://www.whsad.org/wp-content/uploads/2015/01/circleEquation.png Just plug in numbers\[(x-(-7))^2+(y-5)^2 = 3^2\]and simplify

OpenStudy (emsxx3_):

so thats the answer?

OpenStudy (agent0smith):

"simplify" means you need to simplify

OpenStudy (emsxx3_):

but the question don't ask to simplify.

OpenStudy (agent0smith):

Questions never ask you to simplify. They always expect you to know to do that.

OpenStudy (agent0smith):

All you have to do is simplify the -(-7) and 3^2

OpenStudy (peachesncream8):

Hi, the equation of a circle is (x-h)^2 + (y-k)^2 = r^2 . the question has given you the following information the center which is: (-7, 5) also known as the (h,k) of the equation and the radius which is: 3 and r^2 in the question. so all you have to do is : plug in the numbers into the equation: (x+7)^2+ (y-5)^2 = 9 is your answer